OpenStudy (anonymous):

Slove x2+9x+14=0 by using factoring formula

jagr2713 (jagr2713):

Hey do you know how to factor?

OpenStudy (anonymous):

No I don't I forgot :(

OpenStudy (anonymous):

Can you explain me it

jagr2713 (jagr2713):

So we look at the last number, what numbers go into 14 and can be added to 9? because the + is before the 14 makes it + if it was - then a number subtracted that equals to 14

OpenStudy (anonymous):

7 can go into 14?

jagr2713 (jagr2713):

can it be added to 9?

OpenStudy (anonymous):

No I don't think so

jagr2713 (jagr2713):

7*2=14 correct? 7+2?

OpenStudy (anonymous):

=9

jagr2713 (jagr2713):

Great so when it + and + its the same so how would we put this?

jagr2713 (jagr2713):

so we break down x2 into (x+_)(x+_)

jagr2713 (jagr2713):

(x+_)(x+_)=0

OpenStudy (mathmale):

Here's what jagr said earllier: "So we look at the last number, what numbers go into 14 and can be added to 9." To clarify this: Identify FACTORS of the last coefficient (14) and lilst them: 1, 2, 7, 14. Next, choose from this set of numbers the two that ADD UP to 9.

OpenStudy (mathmale):

Note that the 14 is called "the constant term," and that that 9 "the coefficient of the middle term."

jagr2713 (jagr2713):

Yes, when we have a neg and pos, the number next to the last number will be NEG because + means same and - means different

OpenStudy (anonymous):

X+7 and x+3

jagr2713 (jagr2713):

Where did you get 3?

OpenStudy (anonymous):

X+7 and x+9?

jagr2713 (jagr2713):

No, ok lets start over, Ok we look at the last number which is 14. So what number that goes into 14 that CAN be added to get 9. So how 9 is the second number, we are trying to get from 14

OpenStudy (anonymous):

Is it 2 because 7*2=14 and 7+2=9

jagr2713 (jagr2713):

Correct so what do we get?

OpenStudy (mathmale):

Your reasoning is good, appropriate, but what are you asking in "Is it 2?" What are the two binomial factors? Have you checked your work by multiplying these factors together?

jagr2713 (jagr2713):

So we get (X+7)(x+2)=0 When we have the "=0" all we do is put it = to 0 like this x+7=0 and x+2=0 solve :D

OpenStudy (mathmale):

therefore, the two "solutions" of the given equation are x= ? and x= ?

OpenStudy (mathmale):

check these in the original equation, to be sure.

OpenStudy (anonymous):

So x= -7 and X= -2

OpenStudy (mathmale):

Yes. If you substitute these back into your original equation, one at a time, your equation should be TRUE. Nice work!

OpenStudy (anonymous):

Thank you very much :)

OpenStudy (mathmale):

My great pleasure. Appreciate your interest and desire to do well!

OpenStudy (anonymous):

Can you help me on 1 more pls?

OpenStudy (anonymous):

x2-8x=33